### v2

@@ 4.2 Experimental results
To substantiate the efficacy of our technique, we compared our method against established image-to-image conversion methods (Pix2pix [14], Pix2pixHD [15], CycleGAN [22], SPADE [77]), map generation techniques grounded in remote sensing imagery (SMAPGAN [12], Creative GAN [13]), and the current state-of-the-art map generation method, LACG [17].
@@ 4.2.1 Quantitative evaluation
We employed the MLMG to evaluate these models individually, with the results presented in Table 2. Regarding the FID and PSNR metrics, our method consistently achieved optimal outcomes on average. Specifically, on the MLMG-CN dataset, the proposed method recorded an FID index of 103.99 and a PSNR index of 29.759, outperforming other methods. Similarly, the MLMG-US dataset achieved an FID index of 104.63 and a PSNR index of 28.012, demonstrating superior performance. These results indicate that our method surpasses others in these image quality evaluation metrics, effectively validating its efficacy. Notably, the FID results highlight a significant advantage, suggesting that the distribution difference between the maps generated by our method and the real maps is minimal, thus reflecting high-quality output.
Furthermore, we identified suboptimal results in the testing outcomes; for instance, the LCAG method exhibited a suboptimal performance in the FID metric, underscoring the substantial advantage of our method in terms of perceived quality compared to LCAG. In the context of the PSNR metric, our method outperformed the suboptimal SMAPGAN method, indicating higher similarity in the generated maps. However, it is important to note that some researchers have pointed out that a higher PSNR does not necessarily correlate with improved visual quality [78, 79]. We opted to retain this metric as it provides a relative measure of the generated map quality across different methods and is a valuable reference in ablation studies.
We attribute the discrepancies observed in test results between the MLMG-CN and MLMG-US datasets to the inherent differences in the datasets themselves. The MLMG-CN dataset is derived from map data of Chinese cities, while the MLMG-US dataset is based on American city map data. The geographic information within these two datasets varies significantly, resulting in differing test outcomes. Notably, the map data from American cities may exhibit more complex geographic features, whereas the ground features in Chinese cities tend to be more regular, contributing to the lower performance observed in the MLMG-US dataset compared to the MLMG-CN dataset.
Table 2: Quantitative comparison of disparate
methodologies on the MLMG dataset.
Model MLMG-CN MLMG-UN
@@ 4.2.2 Qualitative evaluation
We conducted a visual comparison with all comparative models. Fig. 7 shows that most methods have difficulty generating accurate and continuous cartographic features, such as roads, buildings, and vegetation. We find that multi-scale results from single-scale methods exhibit more ground object errors, feature breaks, and discontinuous tile noise, particularly at higher levels. Map generation methods designed for multi-scale map generation, such as SMAPGAN, CreativeGAN, and the SOTA method LCAG, perform well in multi-scale scale expression. However, their tile results have obvious visual discontinuity problems after map splicing, and the features of the objects they express also have fractures and noise. Compared with other methods, our proposed method is significantly better in terms of the integrity of geographical features and the visual continuity of map tiles. Our method is closer to the ground truth (GT) than other methods, especially in the high-level tile map generation results.
Vertical comparison of the generation outcomes at different levels reveals that other methods produce similar ground object representations, highlighting a problem of inconsistency between map synthesis and map scale. In contrast, the maps generated by our proposed method demonstrate superior synthesis features across various levels. For instance, in the 18-level and 16-level maps, our method showcases more pronounced differences in ground object representation compared to other methods, effectively conveying more accurate map synthesis scale information. These results underscore that our method significantly outperforms the comparison methods in generating multi-scale maps from remote sensing imagery, thereby validating the effectiveness of our approach.
Table 3: Comparison results of two groups of indicators of different model generation effects
@@
C2GM 31.60 0.7503 24.79 0.7351
4.2.3 Map features evaluation
To comprehensively assess the interpretability and spatial consistency of C2GM, we conducted a map feature evaluation using the MFP metric (Section 3.5) on the CSCMG dataset. The results are shown in Table 3. The MFP metric is specifically designed to evaluate the semantic alignment and spatial consistency of generated maps, offering a more comprehensive assessment than traditional metrics. C2GM achieves an MFP score of 0.7351, surpassing TSIT’s score of 0.5910, which indicates enhanced semantic alignment and spatial consistency. The metric’s two components—global feature similarity ($\mathcal{L}_{G}$) and spatial correlation ($\mathcal{L}_{S}$)—effectively capture C2GM’s ability to preserve cartographic semantics, such as road networks and building distributions, and geometric relationships.
C2GM also excels in traditional metrics such as FID (31.60 compared to TSIT’s 109.25) and SSIM (0.7503 compared to 0.7142), confirming its structural coherence and alignment with real map distributions. Although the PSNR values are similar (24.79 compared to 24.87), the MFP metric underscores C2GM’s semantic advantages that traditional feature-level metrics do not capture. Specifically, the sensitivity of the MFP to spatial-topological errors, quantified through self-attention key correlations, provides critical insights beyond those offered by traditional metrics. It establishes C2GM as a robust solution for high-fidelity map generation, supported by both quantitative rigor and qualitative plausibility.
@@ 4.3 Ablation studies
The ablation experiments are divided into two parts. Experiment 1 investigates the impact of incorporating scale information (CLIPEncoder) into the Baseline, illustrating how scale information guidance affects the generated map. Experiment 2 examines the effect of adding cascade references (MFEncoder) to the Baseline, emphasizing the role of tile cascade generation in the map output. Results from these ablation experiments, displayed in Table 4 and Fig. 8, compare the generation outcomes of the Baseline method with those of the proposed method under 2X and 4X cascade references for map denoising. These experiments validate the effectiveness of each component in our approach and identify a more effective span for cascade generation.
Figure 8: Visual results of the ablation study on the CSCMG dataset are presented. Results of our method are visualized after incrementally adding modules. These results demonstrate that the proposed method can generate high-quality and continuous tile map with the assistance of the cascade reference (MFEncoder) and scale information reference (CLIPEncoder).
Table 4: Ablation results for Cascade Reference (MFEncoder) and Scale Information (CLIPEncoder) on the CSCMG dataset.
@@
$\times$ $\times$ 19.962 0.873 52.543
$\times$ 20.898 0.882 44.217
2X $\times$ 23.096 0.934 45.512
2X 31.069 0.941 50.052
4X $\times$ 24.886 0.915 44.329
4X 31.584 0.935 34.369
4.3.1 Ablation study of cascade feature encoder
Adding cascade references to the multi-branch cartographic feature encoder (MFEncoder) aims to fuse low-level map information, further ensuring the consistency of the main content. As indicated in Table 4, FID scores for generated maps with 2X and 4X cascade references are 45.512 and 44.329, respectively, while PSNR scores are 23.096 and 24.886. Compared to the scenario without cascade references, these FID scores reflect increases of 7.031 and 8.241, and PSNR scores show improvements of 3.134 and 4.924, respectively. Such results demonstrate that cascade generation significantly enhances the quality of the generated maps. Furthermore, an improved SSIM index indicates enhanced structural similarity of the generated maps, leading to more accurate outputs.
From Figures 8 (a), (c) and (e), it is evident that considering multi-scale information mitigates misjudgments of geographic elements that arise when only current-level data is analyzed. Visual discontinuities of geographic elements in the original map have been corrected, bringing the generated map closer to the actual representation. Results across different levels illustrate that employing a multi-branch cartographic feature encoder with cascade references allows for integrating single-level generated maps with information from other levels. This integration enables the cartographic model to better recognize objects’ comprehensive features across various scales, thus acquiring more extensive cartographic expertise.
@@ 4.3.1 Ablation study of cascade feature encoder
4.3.2 Ablation study of scale feature adapter
The scale feature adapter is designed to utilize scale information, enabling the generator to optimize output mapping at each scale information and accurately represent content differences. In a map with a scale of 1:8000, the scale feature adapter effectively removes erroneous elements such as an inappropriate yellow background and nonexistent roads. As shown in Table 4, incorporating the scale feature adapter results in improvements across all metrics compared to the baseline, particularly an 8.3 increase in the FID score, highlighting the significant impact of scale information guidance on accurately representing map features. The CLIP-based scale information encoder translates the textual scale data of the generated map into image features, integrates these with ground features extracted from remote sensing images, and embeds them into the pre-trained generator. This process enables the model to discern comprehensive differences among maps of varying scales, leading to improved generation outcomes at each scale that more accurately reflect the characteristics specific to those scales.
@@ 4.3.2 Ablation study of scale feature adapter
Figure 9: Continuous Range Tile Map Generation: C2GM demonstrates powerful capabilities on generating wide-ranging, multi-scale, unbounded continuous, and virtually limitless tile map.
@@ 5 Discussion
5.1 Comparison of the map tile continuous of cascade generated
Notably, the large-scale tile map generation results achieved through the cascading generation method yield high-quality map images. When spliced together, several high-quality generated map tiles form a continuous tile map (Fig. 9). Generative models are highly sensitive to the resolution of input and output images. Leveraging higher resolutions to generate large-scale maps can detrimentally affect generative model performance. A common approach involves stitching together small-scale map tiles to form a larger map [16, 17]. However, this technique often results in visual discontinuities between the tiles, particularly in urban areas where building, road, and other geographic feature boundaries can exhibit noticeable breaks and inconsistencies [12]. This phenomenon is especially pronounced across different map tiles, as they may result from varying generation processes or possess diverse feature distributions. To mitigate this issue, we have adopted a cascaded generation method for producing large-scale maps.
Although we employed a tile-by-tile generation approach, it is evident that the proposed method produces continuous surface features and maintains stylistic similarities between tiles (Fig. 9). Our findings demonstrate that the cascading generation strategy reduces the visual discontinuity typically associated with tile splicing, making large-scale and multi-scale map generation easier. Furthermore, as illustrated in Fig. 7(k), the different scale maps generated for the same area exhibit varying degrees of comprehensiveness; for instance, the 18-level generated maps preserve more intricate details compared to the 16-level maps, much like authentic maps.
5.2 Comparison of the feature generalization of multi-scale generated
The experimental results of multi-scale map generation indicate that using low-level cascade reference map tiles in model training enhances the model’s generalization capability at varying scales (Fig. 9). Consequently, the proposed model can use these generated low-level map tiles as input to produce tile maps with higher resolutions and richer details. Figure 8(f) at map level 16-18 illustrates the multi-scale cascade generation results of C2GM on the CSCMG dataset. Initially, the model is conditioned on existing map tiles at a scale of 1:16000, ultimately generating maps with scales ranging from 1:16000 to 1:2000 based on remote sensing images and cascade references. In contrast to methods that use fixed map layers as constraints, C2GM more effectively captures geographical feature differences between layers through a cascaded reference layers approach, resulting in higher-quality map generation [16, 17]. This distinctive approach facilitates the creation of high-quality, multi-scale, large-scale tile maps that encompass diverse geographical features and scenes.
As a foundation generation model, C2GM unveils new avenues for constructing generative world models within geographic contexts by simulating the cartographic generalization process employed by cartographers during the multi-scale mapping endeavor. The model’s ability to generate maps at different scales, such as 1:2000 and 1:8000, demonstrates its potential for producing high-quality maps with varying levels of detail (Fig. 9). The generated maps exhibit a high degree of realism and continuity, showcasing the model’s effectiveness in capturing the intricate features of the input remote sensing images. This capability is particularly valuable for applications in urban planning, environmental monitoring, and disaster response, where accurate and detailed maps are essential.
@@ 5.3 Limitations and environmental adaptability
The proposed framework demonstrates distinct advantages in urban cartographic scenarios with significant multi-scale generalization requirements (levels 14–18), where artificial features (e.g., buildings, roads) necessitate explicit cartographic operations like aggregation and simplification. However, limitations emerge in natural landscapes and small-scale mapping (levels 1–12). For natural features (e.g., forests, terrain), scale transitions often involve gradual geometric smoothing rather than structural reorganization, reducing the efficacy of cascade learning due to inherent self-similarity across scales. Moreover, tile-based generation relies on localized spatial coherence—a premise valid for structured urban layouts but less so for expansive natural terrains with long-range dependencies.
At small scales (levels 1–12), extreme geometric abstraction transforms features into symbolic representations, creating a resolution gap that challenges pixel-level cascade learning. The current framework’s dependency on explicit feature discrepancies limits its capacity to address such abstraction. While CLIP-based scale encoding partially mitigates global inconsistencies, terrain-aware attention mechanisms remain absent, hindering geomorphological coherence. Future work should explore hybrid vector-raster paradigms to bridge symbolic-geometric disparities and integrate topographic constraints for enhanced natural scene modeling. These adaptations would extend the framework’s applicability while preserving its strengths in urban multi-scale cartography.
@@ 6 Conclusion
This paper introduces C2GM, an innovative foundational framework for generative cartography designed to create multi-scale tile map from remote sensing images. C2GM produces high-quality tile map characterized by rich geographical diversity, seamless visual continuity, and multi-scale representations by enhancing existing generative models. It pioneers multi-scale map synthesis through three synergistic innovations: CLIP-guided multi-scale cascade generation, which integrates cartographic expertise into the diffusion process; self-cascading tile generation with cross-tile consistency constraints to resolve edge discontinuities; and dual-branch scale adaptation for geographically-aware feature extraction. Additionally, we developed the CSCMG dataset, which incorporates cascaded reference and scale information to support the training and evaluation of multi-scale map generation models. The quality of the model’s map generation was evaluated using the MFP metric through feature-level perception. Comprehensive experiments demonstrated C2GM’s exceptional capability in generating multi-scale tile maps.

### v3

@@ 4.2 Experimental results
To substantiate the efficacy of our technique, we compared our method against established image-to-image translation methods (Pix2pix (Isola et al., 2017), Pix2pixHD (Wang et al., 2018), CycleGAN (Zhu et al., 2017), SPADE (Park et al., 2019), SelectionGAN (Tang et al., 2019)), map generation methods (SMAPGAN (Chen et al., 2021), Creative GAN (Fu et al., 2022)), and the current SOTA method, LACG (Fu et al., 2023).
@@ 4.2.1 Quantitative evaluation
SCGM consistently achieved optimal outcomes on average across MLMG datasets, as shown in Table 2. Specifically, on the MLMG-CN dataset, the proposed method recorded an FID index of 103.99 and a PSNR index of 29.759, outperforming other methods. Similarly, the MLMG-US dataset achieved an FID index of 104.63 and a PSNR index of 28.012, demonstrating superior performance. These results indicate that our method surpasses others in these image quality evaluation metrics, effectively validating its efficacy. Notably, the FID results highlight a significant advantage, suggesting that the distribution difference between the maps generated by our method and the real maps is minimal, thus reflecting high-quality output.
Furthermore, we identified suboptimal results in the testing outcomes; for instance, the LCAG method exhibited a suboptimal performance in the FID metric, underscoring the substantial advantage of our method in terms of perceived quality compared to LCAG. In the context of the PSNR metric, our method outperformed the suboptimal SMAPGAN method, indicating higher similarity in the generated maps. However, it is important to note that some researchers have pointed out that a higher PSNR does not necessarily correlate with improved visual quality (Ledig et al., 2017; Pei et al., 2018). We opted to retain this metric as it provides a relative measure of the generated map quality across different methods and is a valuable reference in ablation studies.
Table 2: Quantitative comparison of disparate methodologies on the MLMG dataset, with bold entries indicating optimal results and the underlined entries denoting the suboptimal results.
Model MLMG-CN MLMG-US
@@ 4.2.2 Qualitative evaluation
Visual comparisons (Figure 7) reveal significant limitations in existing methods. Due to tile-edge artifacts and inconsistent object granularity, single-scale approaches fail to maintain coherent geographic features, such as fragmented roads and discontinuous vegetation patterns, particularly at small scales (1:4,000 and 1:2,000, in Figure 7(b) and (c)). Although multi-scale methods, such as SMAPGAN and LCAG, improve hierarchical representation, their tile results have obvious visual discontinuity problems after map splicing, and the features of the objects they express also have fractures and noise, thus undermining cartographic utility (Figure 7(d)). In contrast, SCGM achieves seamless spatial continuity and scale-aware fidelity. For instance, maps at a scale of 1:2,000 preserve intricate urban details, such as building footprints and road networks, while those at 1:8,000 generalize structures without compromising geometric accuracy (Figure 7(e)).
Cross-scale comparison of the generation outcomes at different scales reveals that other methods produce similar ground object representations, represent a problem of inconsistency between map synthesis and map scale. By contrast, our method demonstrate superior synthesis features across various scales: generated tile maps closely align with GT (ground truth) in terms of local coherence (e.g., unbroken linear features) and global consistency (e.g., simplification of geographical features), validating its ability to harmonize multi-scale semantics with generative mapping process. These results underscore that our method significantly outperforms the comparison methods in generating multi-scale maps from remote sensing imagery, thereby validating the effectiveness of our approach.
4.2.3 Map features evaluation
Table 3: Map features evaluation results on the CSCMG dataset.
@@ 4.2.3 Map features evaluation
SCGM 31.60 0.7503 24.79 0.7351
To thoroughly evaluate the interpretability and spatial consistency of SCGM, we conducted a map feature assessment using the MFP metric (Section 3.7) on the CSCMG dataset. Our comparative analysis extended to include representative generative models across two paradigms: GANs, such as Pix2pix, CycleGAN, SMAGAN, and TSIT (Jiang et al., 2020), as well as DMs, including ATME and our proposed SCGM. This comprehensive comparison validates the discriminative power of the MFP metric while highlighting the strengths of SCGM. All models were trained on identical datasets and used the same strategies to ensure fairness. Table 3 summarizes the quantitative results for 1,000 generated map tiles.
The MFP metric—comprising global feature similarity ($\mathcal{L}_{G}$) and spatial correlation ($\mathcal{L}_{S}$)—provides critical insights into semantic alignment and structural fidelity (Section 3.7). SCGM achieved the highest MFP score of 0.7351, significantly outperforming all baselines. Notably, it surpassed the diffusion-based ATME by 11.4%, demonstrating superior preservation of cartographic semantics, such as road topology and building distributions. While GAN-based models exhibited pixel-level similarities (e.g., Pix2pix and CycleGAN attained comparable PSNR values of 23.29 and 24.10, respectively), their low MFP scores (0.4834 and 0.5357) reflect inherent limitations in spatial-semantic consistency. This disparity underscores MFP’s sensitivity to semantic discrepancies that conventional pixel-wise metrics fail to capture. For instance, SMAGAN achieved competitive SSIM (0.8204) and PSNR (24.29) scores but registered the lowest MFP (0.3538), indicating that it produced structurally plausible yet semantically inconsistent outputs. In contrast, SCGM’s balanced excellence across all metrics, including a leading FID score of 31.60 and an SSIM score of 0.7503, validates its comprehensive generation capability.
Traditional metrics further support these findings. The substantial FID gap between SCGM (31.60) and TSIT (109.25) highlights SCGM’s alignment with real map distributions. While SCGM’s PSNR (24.79) is slightly lower than TSIT’s (24.87), this marginal difference highlights that pixel-level fidelity alone cannot guarantee cartographic validity—a limitation addressed by MFP’s emphasis on spatial-topological relationships. The metric’s quantification of self-attention key correlations in spatial errors provides critical insights into model behavior. Our analysis establishes that SCGM advances the state-of-the-art in multi-scale map generation by coordinating map scale modalities with feature representation. It positions SCGM as a robust solution for high-fidelity map generation, supported by both quantitative rigor and qualitative plausibility.
@@ 4.3 Ablation studies
In this section, we perform ablation studies on two core components of SCGM: the cascading feature encoder and the scale-modality encoder, as well as on cascade spans (2X/4X), to validate their necessity and optimal configurations. All experiments are conducted on the CSCMG dataset.
Figure 8: Incremental visual results showcasing seamless, high-quality tile map generation with the addition of cascade referencing (MFEncoder) and scale encoding (ScaleEncoder). The orange box highlights fine-grained map details contributed by cascade references, while the red box denotes enhanced scale representation from scale encoding.
Table 4: Ablation results for MFEncoder and ScaleEncoder on CSCMG dataset, where "w/ 2X" and "w/ 4X" indicate conditional generation using 2X or 4X cascade references, respectively.
@@
w/o w/o 19.962 0.873 52.543
w/o w/ 20.898 0.882 44.217
w/ 2X w/o 23.096 0.934 45.512
w/ 2X w/ 31.069 0.941 50.052
w/ 4X w/o 24.886 0.915 44.329
w/ 4X w/ 31.584 0.935 34.369
4.3.1 Effects of MFEncoder
Integrating cascade references into the MFEncoder significantly enhances map generation quality, as evidenced by quantitative metrics (Table 4). FID scores for generated maps with 2X and 4X cascade references are 45.512 and 44.329, respectively, while PSNR scores are 23.096 and 24.886. Compared to the scenario without cascade references, these FID scores reflect increases of 7.031 and 8.241, and PSNR scores show improvements of 3.134 and 4.924, respectively. Such results demonstrate that cascade generation significantly enhances the quality of the generated maps. Furthermore, an improved SSIM index indicates enhanced structural similarity of the generated maps, leading to more accurate outputs.
From Figures 8 (a), (c) and (e), it is evident that considering multi-scale information mitigates misjudgments of geographic elements that arise when only current-scale data is analyzed. Visual discontinuities of geographic elements in the original map have been corrected, bringing the generated map closer to the actual representation. Results across different scales illustrate that employing a multi-branch cartographic feature encoder with cascade references allows for integrating single-scale generated maps with information from other scales. This integration enables the mapping model to better recognize objects’ comprehensive features across various scales, thus acquiring more extensive cartographic expertise.
@@ 4.3.1 Effects of MFEncoder
4.3.2 Effects of ScaleEncoder
The ScaleEncoder ensures an accurate representation of scale-specific cartographic semantics. In a map with a scale of 1:8000, the ScaleEncoder effectively removes erroneous elements such as an inappropriate yellow background and nonexistent roads. As shown in Table 4, incorporating the ScaleEncoder results in improvements across all metrics compared to the baseline, particularly an 8.3 increase in the FID score, highlighting the significant impact of scale information guidance on accurately representing map features. The CLIP-based scale information encoder translates the textual scale data of the generated map into image features, integrates these with ground features extracted from remote sensing images, and embeds them into the pre-trained generator. This process enables the model to discern comprehensive differences among maps of varying scales, leading to improved generation outcomes at each scale that more accurately reflect the characteristics specific to those scales.
@@ 5 Discussion
Figure 9: Seamless tile map generation: SCGM demonstrates powerful capabilities on generating seamless, multi-scale, large-format, and virtually limitless tile maps.
5.1 Analysis of multi-scale cascade generation
The integration of cascade references enhances multi-scale feature generalization. By leveraging small-scale tiles (e.g., 1:15,000 scale) as priors, SCGM generates higher-scale outputs (e.g., 1:8000 to 1:2,000) with enriched geographic semantics (Figure 8(f)). This cascaded refinement captures scale-specific feature hierarchies—such as road network granularity and vegetation patterns—more effectively than fixed-layer constraints (Li et al., 2020; Fu et al., 2022). The framework’s ability to simulate cartographic generalization processes enables the synthesis of realistic, seamless tile maps across diverse scales (Figure 9), supporting applications like disaster response and urban planning where accuracy and detail adaptability are paramount.
Furthermore, the cascading generation strategy in SCGM addresses critical challenges in large-format multi-scale map synthesis. Unlike conventional tile stitching methods (Liu et al., 2021; Fu et al., 2023), which introduce visual discontinuities at feature boundaries (e.g., roads, buildings), SCGM enforces spatial coherence by conditioning high-level tile generation on lower-scale outputs (Figure 9). This approach mitigates edge artifacts common in independent tile generation (Chen et al., 2021), ensuring stylistic consistency and geographic continuity across scales. For instance, urban maps at 1:2,000 scale retain intricate details (e.g., building footprints), while 1:8,000 scale maps exhibit generalized structures, mirroring cartographic conventions (Figure 7(k)).
@@ 5.2 Limitations and environmental adaptability
While SCGM excels in urban mapping (scales 1:8,000 to 1:2,000) by handling artificial features (e.g., buildings, roads) through aggregation and simplification, its performance declines in natural landscapes and small-scale mapping (<1:35,000). For natural features (e.g., forests, terrain), scale transitions rely on gradual geometric smoothing rather than structural reorganization, limiting cascaded learning efficacy due to inherent self-similarity. Tile-based generation assumes localized spatial coherence, which suits urban layouts but fails for expansive terrains with long-range dependencies. At small scales (<1:35,000), geometric abstraction converts features into symbolic representations, challenging pixel-level cascade learning. Although CLIP-based scale encoding mitigates global inconsistencies, the absence of terrain-aware attention mechanisms hinders geomorphological fidelity. Future work should focus on enhancing natural scene modeling while preserving urban mapping strengths.
@@ 6 Conclusion
This paper introduces SCGM, an innovative foundational framework for generative mapping designed to create multi-scale tile map from remote sensing images. The primary contributions of this study consist of the following: First, a CLIP-guided scale-modality encoding mechanism that dynamically aligns cartographic generalization rules (e.g., resolution, legends) with geographic semantics. Second, a cascaded generation strategy that enforces cross-tile consistency by conditioning larger-scale outputs on smaller-scale priors, eliminating tile discontinuities. Third, a dual-branch adaptation network that hierarchically fuses multi-scale features while preserving geographic fidelity. When evaluated on the CSCMG dataset, SCGM achieves state-of-the-art performance in multi-scale tile map generation. It ensures seamless continuity and incorporates scale-aware cartographic features, as measured by the map feature perception metric. The framework’s capability to simulate cartographic workflows opens new avenues for real-time applications in emergency mapping and urban planning.
Data availability
The SCGM framework and CSCMG dataset for our key contributions will be open sourced in the code repository: https://github.com/Magician-MO/SCGM.